Given a sample of 100 projector bulbs from a company has a mean

length of life of 20.5 hours with a standard deviation of 1.6 hours, how do I find a 95% confidence interval for the average length of life of those bulbs and then interpret the results?


1
Expert's answer
2022-04-05T07:37:30-0400

The formula to calculate a confidence interval for a population mean is as follows:

"CI=\bar{x}\pm z\cdot\cfrac{\sigma}{\sqrt{n}},"

where:

  • "\bar{x}=20.5" : sample mean
  • zthe chosen z-value, for a 95% confidence interval z = 1.96
  • "\sigma=1.6"sample standard deviation
  • n = 100 sample size.

So,

"CI=20.5 \pm 1.96\cdot\cfrac{1.6}{\sqrt{100}}=20.5\pm0.314=(20.186, 20.814)."


There is a 95% chance that the confidence interval of (20.186, 20.814) contains the true population mean length of life of bulbs.

Another way of saying the same thing is that there is only a 5% chance that the true population mean lies outside of the 95% confidence interval. That is, there’s only a 5% chance that the true population mean length of life of bulbs is greater than 20.814 hours or less than 20.186 hours.